# A Survey on Adversarial Contention Resolution

### v1

@@ 3.1. AQT and Contention Resolution
Another important metric is (packet) latency, which is often defined as the maximum time that any packet stays in the queue prior to being successfully transmitted. We note that, while this is described in the context of queues, this corresponds to the metric as described in Section 1.1.3.
@@
Given the prominence of backoff algorithms, one may wonder about their performance in the AQT setting. Given the suboptimal makespan of popular algorithms like BEB in the static setting, we should be skeptical of their performance when packets arrive dynamically at a sufficiently high rate. To help solidify this intuition, we sketch an argument that shows the instability of an algorithm BEB’, which is very similar to BEB. Specifically, for each slot $i\geq 1$ starting with the slot that packet $u$ injected into the system, $u$ sends with probability $1/i$. Note that this is similar to the Bernoulli-backoff algorithm (recall Section 2.1), where the sending probability per slot proceeds as $1/2,1/2,1/4,1/4,1/4,1/4,1/8,1/8,...$ etc. By comparison, BEB’ proceeds as $1,1/2,1/3,1/4,1/5,1/6,...$, and so at any given slot since a packet is injected, the sending probabilities under the Bernoulli-backoff algorithm and BEB’ differ by only a positive constant. [^5] We highlight that this is very similar to the algorithm Decrease Slowly in (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ); see the proof of Theorem 8 for a formal analysis.
@@
Overview of Literature. As mentioned above, the result by Bender et al. (bender:adversarial, ) explores stability of backoff algorithms under windowed adversaries in the queue-free model. BEB is shown to be stable for $\rho=d/\log n$ for some constant $d>0$, and unstable for $\rho=\Omega(\log\log n/\log n)$ They also show that LLB is stable for $\rho=d^{\prime}/((\log n)\log\log n)$ for some constant $d^{\prime}>0$, and unstable for $\rho=\Omega(1/\log n)$. Overall, traditional backoff algorithms appear to perform sub-optimally, although this is not surprising given the corresponding makespan results for the static setting.
@@
Additional impossibility results are made in the context of three natural properties. First, an algorithm is fair if each packet is transmitted successfully within $O(b/\rho)$ slots from the time it is injected. The authors show that no algorithm can be fair and stable for any number of at least $2$ participants against a leaky-bucket adversary of burstiness $2$. Second, an algorithm is withholding the channel if, upon successfully transmitting a packet, the station keeps sending until all of its pending packets are exhausted. No algorithm that withholds the channel can be stable for three stations against a leaky-bucket adversary of burstiness $2$. Third, an algorithm is said to be retaining if after a participant successfully transmits a packet, and if any packets arriving subsequent to this event are injected only at other stations, the participant eventually empties its queue. No retaining protocol is stable in a system of at least four stations against a window adversary of burstiness $2$. Chlebus et al. show that this implies no protocol that is fair can also be stable in a system of at least four stations against the window adversary of burstiness $2$.
@@
Anantharamu et al. (anantharamu2010deterministic, ) and Aldawsari et al. (aldawsari:adversarial, ) design and analyze several deterministic algorithms that employ either a round-robin or token-passing approach for obtaining access to the channel. The authors derive bounds on latency under the leaky-bucket adversary model as a function of $\rho$, for $\rho<1/(2\log n)$ and general $b$. Note that this restriction on $\rho$ is unsurprising, given the negative results on strong stability by Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) and prior results on the challenges posed by burstiness (chlebus2009maximum, ; chlebus2007stability, ). These algorithms are compared, both analytically and via simulations, against the randomized algorithms, BEB and Quadratic Backoff (QB), where the latter has each successive window increase as the square of the previous window. The simulation results illustrate the impact of different injection rates, burstiness values, and the number of participants. These findings generally align with the theoretical predictions and, interestingly, the deterministic algorithms compare favorably to BEB and QB, suggesting that randomized approaches may not necessarily be superior. More-recent work by Hradovich et al. (HradovichKK21, ) provides a new average-case analysis of the maximum queue size and latency for several of the same algorithms.
@@
Anantharamu et al. (anantharamu2009adversarial, ; anantharamu2017adversarial, ) consider deterministic full-sensing and acknowledgment-based algorithms for the case where each participant is subject to an individual injection rate under a window adversary, where the window $w$ is unknown. There are $n$ participants, each with a unique ID in ${1,...,n}$, and $n$ is known. The authors prove a lower bound on latency of $\Omega(w\log n/\log w)$ for $w\leq n$, and $\Omega(w)$ for $w>n$. Upper bounds are established by a full-sensing algorithm with CD, and by an adaptive algorithm without CD, each of which achieve a packet latency of $O(\min(n+w,w\log n))$; Anantharamu et al. also give a full-sensing algorithm without CD that has $O(n+w)$ queue size and $O(nw)$ latency.
In contrast to the above, the remaining results address a model where participants have unique IDs in $\{1,...,n\}$, where $n$ is known. Additionally, participants have access to a global clock. Hradovich et al. (hradovich2020contention, ) consider a problem variant where, in any fixed round, at most $k$ stations may access the multiple access channel; this is called a $k$-restrained channel.The authors design and analyze deterministic algorithms for leaky bucket adversaries with $\rho\leq 1$. The first is a stable adaptive algorithm whose maximum number of queued packets is $O(n^{2}+b)$, which is reminiscent of the bound in (chlebus2009maximum, ). The second is a full-sensing algorithm that is stable for $\rho\leq 1-1/n$. Finally, the authors give an acknowledgment-based algorithm with throughput $\Theta(k/n(\log^{2}n))$, which is complemented by a lower bound that no acknowledgement-based algorithm can be stable for a rate at least $\min\{k/n,1/(3\log n)\}$. Simulations indicate that the proposed algorithms perform favorably, with respect to queue size and number of channel accesses, relative to well-known backoff algorithms.
Bienkowski et al. (bienkowski:distributed, ; BienkowskiJKK12, ) consider a model where each participant has a queue, and messages may include control bits. The adversary considered here differs from the leaky-bucket and window adversaries discussed earlier: at any time, an arbitrary number of packets can be injected into the queue of any participant. The metrics of interest are the total packets in all queues, called total load, and the maximum queue size. The authors prove that their deterministic algorithm achieves a total load that is an additive $O(n^{2})$ amount larger than any other algorithm, and the maximum queue size is at most an $n$-factor larger.
@@
More recently, Garncarek et al. (garncarek2019stable, ) operate in the same model, with a focus on algorithms that keep very little or no information about the previous state of the system. Here, the authors show a deterministic algorithm that is stable for $\rho=\Omega(1/\log n)$ given that only $O(1)$ bits are used as memory to store some history of events, and each participant can check its queue size (not just whether it is empty). Stability can also be achieved when the rate is increased, and with zero bits of memory; however, $b$ must be upper bounded by some value known a priori.
Discussion. There are several significant messages conveyed by the literature surveyed above. Randomized backoff algorithms continue to disappoint in that they are stable only for sub-constant injection rates. Interestingly, in terms of latency, deterministic algorithms can achieve performance that rivals and even surpasses traditional randomized backoff algorithms. However, most of these results make strong assumptions about unique IDs and/or a priori knowledge of $n$. The exception is the algorithm by Anantharamu and Clebus (anantharamu2015broadcasting, ), which assumes a $1$-activating adversary. This leads to a more general message: while stability and fairness seem like natural qualities that we desire in algorithm design, it is often not possible to achieve both simultaneously when there is even a little (i.e., constant) burstiness. Therefore, all of this paints a fairly dismal picture of what is possible under window and leaky bucket adversaries, especially for deterministic algorithms, given the many impossibility results that apply.
What avenues are there for positive results? As we will discuss later in Section 3.3, new randomized approaches can offer better performance. However, fundamentally, perhaps the AQT model is overly pessimistic; this has been investigated in more general settings (e.g., (BKS14, ).) Here, focusing on whether queue sizes can become unbounded under a very specific injection pattern, and under an infinite stream of packets, may lead to results that are artificially bleak. In the next section, we turn our attention to work that is concerned with a finite stream of packets that arrive under an arbitrary schedule, and the makespan is an important performance metric.
3.2. Deterministic Approaches for Contention Resolution
@@
In this section, we depart from parameterized adversaries and notions of stability, while still focusing on deterministic contention resolution algorithms. Here, our performance metric is latency (recall Section 1.1.3). Much of the contention-resolution work in this setting is referred to as $k$-selection or the wakeup problem. The former addresses $n$ participants of which $k$, for $1\leq k\leq n$, activate over time, each with a single packet to transmit. The activation times are selected by an adversary, where stations are operational from the beginning of time, and packets arrive subject to parameters $\rho$ and $b$. As before, most of the literature operates in a model where each participant has a unique ID. Furthermore, $n$ is typically assumed to be known; throughout this section, we assume this to be true unless explicitly stated otherwise. The goal is to minimize the latency (recall from Section 1.1 ) required for all $k$ participants to successfully send their respective packets.
The wakeup problem also considers $n$ participants, of which some are awake and the remainder are inactive (or“sleeping”), but monitoring the channel. Devices may awaken to begin executing an algorithm as scheduled in some worst case fashion, thus potentially causing contention on the shared channel. Each device has a packet to send, and upon the first success, all sleeping devices awaken and the execution ends. [^6] Note that, given that we require only a single success for the wake up problem, it does not make sense to categorize wakeup protocols as acknowledgement or non-acknowledgement-based. Here, minimizing the time until all devices are awake is the goal.
@@
Many results put a spin on John von Neumann’s quote by generating deterministic results from randomized approaches. Specifically, a common technique in this area is to design a randomized algorithm, and then use the probabilistic method (alon2016probabilistic, ) to argue that a deterministic equivalent exists. Here, we focus on the result by De Marco and Kowalski (doi:10.1137/140982763, ); however, related approaches are used in several other papers to give existence proofs for efficient transmission schedules (DBLP:conf/icdcs/MarcoKS19, ; de2013contention, ; DEMARCO20171, ; Chlebus:2016:SWM:2882263.2882514, ; ChlebusK04, ; ChlebusGKR05, ).
De Marco and Kowalski (doi:10.1137/140982763, ) consider the dynamic $k$-selection problem where there is a global clock, $k$ is unknown, and no-CD. Their main result is a non-adaptive, deterministic algorithm for contention resolution with latency $O(k\log(n)\log\log(n))$. Notably, this result almost matches a previous lower bound by Greenberg and Winograd (greenberg1985lower, ) in $\Omega(k\log(n)/\log k)$ for adaptive algorithms under the static case with collision detection. Therefore, this result by De Marco and Kowalski illustrates that it is possible to achieve near-optimal latency with a non-adaptive and deterministic algorithm.
@@
How is membership in $\mathcal{M}_{i,j}$ determined? This is decided randomly: participant $u$ belongs to $\mathcal{M}_{i,j}$ with probability approximately $\frac{1}{2^{i}2^{j\mod\lg\lg n}}$, for $1\leq i\leq\lg n$ and $j\geq 1$. Note that the first term of the sending probability—the $1/2^{i}$ term—and also the interval length—the $2^{i}$ term—are reminiscent of BEB; in particular, they seem analogous to the $i^{\mbox{\tiny th}}$ window of the Bernoulli version (recall Sections 2.1 and 3.1). Consequently, we should not be surprised at a latency that is $\Omega(n\log n)$.
How does contention behave under $\mathcal{M}$? For illustrative purposes, consider $\Theta(k)$ participants executing the schedule in rows $1$ and $\lg k$. In row $1$, an active participant $u$ may contribute significantly to contention, but only for its small interval $\mathcal{I}_{1,u}$. For example, if $\Theta(k)$ participants all activate simultaneously (i.e., a batch), then their contribution to contention is $\Theta(k)$ and no participant succeeds; however, this occurs only over $\Theta(\log n\log\log n)$ slots in row $1$. Conversely, the $\Theta(k)$ participants may be spread out over row $1$, each contributing only $O(1)$ contention in their own respective, disjoint interval of $\Theta(\log n\log\log n)$ slots; we will return to these two extremems momentarily.
@@
Recall the contribution to contention from participants in row $1$ using the two qualitatively different cases discussed above: their contribution may be high for “small” number—specifically, $\Theta(\log n\log\log n)$—slots of $\mathcal{I}_{\lg(k)}$, or only $O(1)$ over perhaps all of $\mathcal{I}_{\lg(k)}$. Therefore, accounting for these two cases, in row 1 there should be many disjoint periods of $\Theta(\log\log n)$ contiguous slots where contention is good. More generally, given the contribution to contention from other rows, the authors show that, w.h.p. in $n$, an important property holds: in every interval of length $\Theta(k\log(n)\log\log(n))$ slots, $O(k/\log n)$ participants proceed from row $\lg(k)$ to row $\lg(k)+1$ of the transmission matrix. De Marco and Kowalski (doi:10.1137/140982763, ) show that a random transmission matrix yields the above property with strictly positive probability. Therefore, by the probabilistic method, the authors argue that there must exist a matrix with this property.
But what about these $O(k/\log n)$ participants that keep executing in row $\lg(k)+1$? Interleaved with the execution is the deterministic result by Komlós and Greenberg (komlos:asymptotically, ) (recall Section 2.2), which addresses the static setting. Specifically, given a batch of $m$ out of $n$ participants with a packet to send, where both $m$ and $n$ are known, all $m$ participants succeed in time $O(m+m\log(n/m))$. Here, in the dynamic $k$-selection problem, $m$ is not known, and so the result by Komlós and Greenberg is executed for values of $m=1,2,...,\lg n$, which contributes an extra logarithmic-in-$n$ factor to the run time. Thus, the $m=O(k/\log n)$ participants that remain after row $k$ succeed within an additional $O(m+m\log(n/m))\log n$, which is $O(k\log(n\log n))$ by substituting in for $m$.

### v3

@@ 3.1. AQT and Contention Resolution
Another important metric is (packet) latency, which is often defined as the maximum time that any packet stays in the queue prior to being successfully transmitted. We note that, while this is described in the context of queues, this aligns with the metric as described in Section 1.1.3.
@@
Given the prominence of backoff algorithms, one may wonder about their performance in the AQT setting. Given the suboptimal makespan of popular algorithms like BEB in the static case, we should be skeptical of their performance when packets arrive dynamically at a high rate. To further this intuition, we sketch an argument that shows the instability of an algorithm BEB’, which is similar to BEB. Specifically, for each slot $i\geq 1$ starting with the slot that packet $u$ injected into the system, $u$ sends with probability $1/i$. Note that this is similar to the Bernoulli-backoff algorithm (recall Section 2.1), where the sending probability per slot proceeds as $1/2,1/2,1/4,1/4,1/4,1/4,1/8,1/8,...$ etc. By comparison, BEB’ proceeds as $1,1/2,1/3,1/4,1/5,...$, and so at any given slot since a packet is injected, the sending probabilities under the Bernoulli-backoff algorithm and BEB’ differ by only a constant. [^6] We highlight that this is very similar to the algorithm Decrease Slowly given by Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ); see the proof of Theorem 8 in (jurdzinski2005probabilistic, ) for a formal analysis.
@@
Overview of Literature. As mentioned above, the result by Bender et al. (bender:adversarial, ) explores stability of backoff algorithms under windowed adversaries in the queue-free model. BEB is shown to be stable for $\rho=d/\log n$ for some constant $d>0$, and unstable for $\rho=\Omega(\log\log n/\log n)$ They also show that LLB is stable for $\rho=d^{\prime}/((\log n)\log\log n)$ for some constant $d^{\prime}>0$, and unstable for $\rho=\Omega(1/\log n)$. Overall, traditional backoff algorithms appear to perform sub-optimally, although this is not surprising given the corresponding makespan results for the static case.
@@
Additional impossibility results are made in the context of three natural properties. First, an algorithm is fair if each packet is transmitted successfully within $O(b/\rho)$ slots from the time it is injected. [^7] Note that this is a different notion of fairness than described in Section 2.3. The authors show that no algorithm can be fair and stable for any number of at least two participants against a leaky-bucket adversary of burstiness $2$. Second, an algorithm is withholding the channel if, upon successfully transmitting a packet, the station keeps sending until all of its pending packets are exhausted. No algorithm that withholds the channel can be stable for three stations against a leaky-bucket adversary of burstiness $2$. Third, an algorithm is said to be retaining if after a participant successfully transmits a packet, and if any packets arriving subsequent to this event are injected only at other stations, the participant eventually empties its queue. No retaining protocol is stable in a system of at least four stations against a window adversary of burstiness $2$. Chlebus et al. show that this implies no protocol that is fair can also be stable in a system of at least four stations against the window adversary of burstiness $2$.
@@
Anantharamu et al. (anantharamu2010deterministic, ) and Aldawsari et al. (aldawsari:adversarial, ) design and analyze several deterministic algorithms that employ either a round-robin or token-passing approach for obtaining access to the channel. The authors derive bounds on latency under the leaky-bucket adversary model as a function of $\rho$, for $\rho<1/(2\log n)$ and general $b$. Note that this restriction on $\rho$ should not surprise us, given the negative results on strong stability by Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) and prior results on the challenges posed by burstiness (chlebus2009maximum, ; chlebus2007stability, ). These algorithms are compared, both analytically and via simulations, against the randomized algorithms, BEB and Quadratic Backoff (QB), where the latter has each successive window size increase as the square of the previous window. The simulation results illustrate the impact of different injection rates, burstiness values, and the number of participants. These findings generally align with the theoretical predictions and, interestingly, the deterministic algorithms compare favorably to BEB and QB, suggesting that randomized approaches may not necessarily be superior. More-recent work by Hradovich et al. (HradovichKK21, ) provides a new average-case analysis of the maximum queue size and latency for several of the same algorithms.
@@
Anantharamu et al. (anantharamu2009adversarial, ; anantharamu2017adversarial, ) consider deterministic full-sensing and acknowledgment-based algorithms for the case where each participant is subject to an individual injection rate under a window adversary, where the window $w$ is unknown. There are $n$ participants, each with a unique ID in ${1,...,n}$, and $n$ is known. The authors prove a lower bound on latency of $\Omega(w\max\{1,\log_{w}n\})$ for $w\leq n$, and $\Omega(w+n)$ for $w>n$. Upper bounds are established by a full-sensing algorithm with CD, and by an adaptive algorithm without CD, each of which achieve a packet latency of $O(\min(n+w,w\log n))$; Anantharamu et al. also give a full-sensing algorithm without CD that has $O(n+w)$ queue size and $O(nw)$ latency.
In contrast to the above, the remaining results address a model where participants have unique IDs in $\{1,...,n\}$, where $n$ is known. Additionally, participants have access to a global clock. Hradovich et al. (hradovich2020contention, ) consider a problem variant where, in any fixed round, at most $k$ stations may access the multiple access channel; this is called a $k$-restrained channel. The authors design and analyze deterministic algorithms for leaky bucket adversaries with $\rho\leq 1$. The first is a stable adaptive algorithm whose maximum number of queued packets is $O(n^{2}+b)$, which is reminiscent of the bound in (chlebus2009maximum, ). The second is a full-sensing algorithm that is stable for $\rho\leq 1-1/n$. Finally, the authors give an acknowledgment-based algorithm with throughput $\Theta(k/n(\log^{2}n))$, which is complemented by a lower bound that no acknowledgement-based algorithm can achieve a throughput better than $\min\{k/n,1/(3\log n)\}$. Simulations indicate that the proposed algorithms perform favorably, with respect to queue size and number of channel accesses, relative to well-known backoff algorithms.
Bienkowski et al. (bienkowski:distributed, ; BienkowskiJKK12, ) consider a model where each participant has a queue, and messages may include control bits. The adversary considered here differs from the leaky-bucket and window adversaries discussed earlier: at any time, an arbitrary number of packets can be injected into the queue of any participant. The metrics of interest are the total packets in all queues, called total load, and the maximum queue size. The authors prove that their deterministic algorithm achieves a total load that is an additive $O(n^{2})$ amount larger than the best offline algorithm, and the maximum queue size is at most an $n$-factor larger.
@@
More recently, Garncarek et al. (garncarek2019stable, ) operate in the same model, with a focus on algorithms that keep very little or no information about the previous state of the system. Here, the authors show a deterministic algorithm that is stable for $\rho=\Omega(1/\log n)$ given that only $O(1)$ bits are used as memory to store some history of events, and each participant can check its queue size (not just whether it is empty). Stability can also be achieved when the rate is increased to a constant, and with zero bits of memory; however, $b$ must be upper bounded by some value known a priori.
Summary. There are several significant messages conveyed by the literature surveyed above. Randomized backoff algorithms continue to disappoint in that they are stable only for sub-constant injection rates. Interestingly, in terms of latency, deterministic algorithms can achieve performance that rivals and even surpasses traditional randomized backoff algorithms. However, most of these results make assumptions about unique IDs and/or a priori knowledge of $n$. The exception is the algorithm by Anantharamu and Clebus (anantharamu2015broadcasting, ), which assumes a $1$-activating adversary. This leads to a more general message: while stability and fairness seem like natural qualities that we desire in algorithm design, it is often not possible to achieve both simultaneously when there is even a little (i.e., constant) burstiness. Therefore, all of this paints a fairly dismal picture of what is possible under window and leaky bucket adversaries, especially for deterministic algorithms, given the many impossibility results that apply.
What avenues are there for positive results? As we will discuss later in Section 3.3, new randomized approaches can offer better performance. However, fundamentally, perhaps the AQT model is overly pessimistic; this has been investigated in more general settings (e.g., (BKS14, ).) Here, focusing on whether queue sizes can become unbounded under a very specific injection pattern, and under an infinite stream of packets, may lead to results that are artificially bleak. In the next section, we turn our attention to work that is concerned with a finite stream of packets that arrive under an arbitrary schedule.
3.2. Deterministic Approaches
@@
In this section, we depart from parameterized adversaries and notions of stability, while still focusing on deterministic contention resolution algorithms. Here, our performance metric is latency (recall Section 1.1.3). Much of the contention-resolution work in this setting is referred to as $k$-selection, which is described as follows. There are $n$ participants of which $k$, for $1\leq k\leq n$, activate over time, each with a single packet to transmit. The activation times are selected by an adversary. As before, most of the literature operates in a model where each participant has a unique ID. Furthermore, $n$ is typically assumed to be known; throughout this section, we assume this to be true unless explicitly stated otherwise. The goal is to minimize the latency (recall from Section 1.1 ) required for all $k$ participants to successfully send their respective packets.
@@
Many results put a spin on John von Neumann’s quote by generating deterministic results from randomized approaches. Specifically, a common technique in this area is to design a randomized algorithm, and then use the probabilistic method (alon2016probabilistic, ) to argue that a deterministic analog exists. Here, we focus on the result by De Marco and Kowalski (doi:10.1137/140982763, ); however, related approaches are used in several other papers to give existence proofs for efficient transmission schedules (DBLP:conf/icdcs/MarcoKS19, ; de2013contention, ; DEMARCO20171, ; Chlebus:2016:SWM:2882263.2882514, ; ChlebusK04, ; ChlebusGKR05, ).
De Marco and Kowalski (doi:10.1137/140982763, ) consider the dynamic $k$-selection problem where there is a global clock, $k$ is unknown, and no-CD. Their main result is a non-adaptive, deterministic algorithm for contention resolution with latency $O(k\log(n)\log\log(n))$. Notably, this result almost matches a previous lower bound by Greenberg and Winograd (greenberg1985lower, ) in $\Omega(k\log(n)/\log k)$ for adaptive algorithms under the static case with collision detection.
@@
How is membership in $\mathcal{M}_{i,j}$ determined? This is decided randomly: participant $u$ belongs to $\mathcal{M}_{i,j}$ with probability approximately $\frac{1}{2^{i}2^{j\mod\lg\lg n}}$, for $1\leq i\leq\lg n$ and $j\geq 1$. Note that the first term of the sending probability (the $1/2^{i}$ term) and also the interval length (specifically, the $2^{i}$ term) are reminiscent of the $i^{\mbox{\tiny th}}$ window for the Bernoulli version of BEB (recall Sections 2.1 and 3.1). Consequently, we should not be surprised at a $\Omega(k\log n)$ latency.
How does contention behave under $\mathcal{M}$? For illustrative purposes, consider $\Theta(k)$ participants executing the schedule in rows $1$ and $\lg k$. In row $1$, an active participant $u$ may contribute significantly to contention, but only for its small interval $\mathcal{I}_{1,u}$. For example, if $\Theta(k)$ participants all activate simultaneously (i.e., a batch), then their contribution to contention is $\Theta(k)$ and no participant succeeds; however, this occurs only over $\Theta(\log n\log\log n)$ slots in row $1$. Conversely, the $\Theta(k)$ participants may be spread out over row $1$, each contributing only $O(1)$ contention in their own respective, disjoint interval of $\Theta(\log n\log\log n)$ slots; we will return to these two extremes momentarily.
@@
Recall the contribution to contention from participants in row $1$ using the two qualitatively different cases discussed above: their contribution may be high for a “small” number—specifically, $\Theta(\log n\log\log n)$—slots of $\mathcal{I}_{\lg(k)}$, or only $O(1)$ over perhaps all of $\mathcal{I}_{\lg(k)}$. Therefore, accounting for these two cases, in row 1 there should be many disjoint periods of $\Theta(\log\log n)$ contiguous slots where contention is good. More generally, given the contribution to contention from other rows, the authors show that, w.h.p. in $n$, an important property holds: in every interval of length $\Theta(k\log(n)\log\log(n))$ slots, $O(k/\log n)$ participants proceed from row $\lg(k)$ to row $\lg(k)+1$ of the transmission matrix. De Marco and Kowalski (doi:10.1137/140982763, ) show that a random transmission matrix yields the above property with strictly positive probability. Therefore, by the probabilistic method, the authors argue that there must exist a matrix with this property.
But what about these $O(k/\log n)$ participants that keep executing in row $\lg(k)+1$? Interleaved with the execution is the deterministic result by Komlós and Greenberg (komlos:asymptotically, ) (recall Section 2.2), which addresses the static case. Specifically, given a batch of $m$ out of $n$ participants with a packet to send, where both $m$ and $n$ are known, all $m$ participants succeed in time $O(m+m\log(n/m))$. Here, in the dynamic $k$-selection problem, $m$ is not known, and so the result by Komlós and Greenberg is executed for values of $m=1,2,...,\lg n$, which contributes an extra logarithmic-in-$n$ factor to the run time. Thus, the $m=O(k/\log n)$ participants that remain after row $k$ succeed within an additional $O(m+m\log(n/m))\log n$, which is $O(k\log(n\log n))$ by substituting in for $m$.